If u = x2 + y2
then the value of
1
Expert's answer
2020-11-09T14:59:46-0500

"\displaystyle

u(x, y) = x^2 + y^2\\


\frac{\partial u}{\partial x} = 2x, \frac{\partial^2 u}{\partial x^2} = 2\\


\frac{\partial u}{\partial y} = 2y, \frac{\partial^2 u}{\partial y^2} = 2\\


\textsf{Since}\, \frac{\partial^2 u}{\partial x^2} + \frac{\partial^2 u}{\partial y^2} = 4\,,u\, \textsf{is not harmonic.}\\

u\, \textsf{is not regular and has singularities}\\
\textsf{at every point except at}\,x = y = 0."